1. Passwords are made using 4 letters and 2 digits. How many passwords can be made if repetition of letters and digits is allowed?
Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

1. Passwords are made using 4 letters and 2 digits. How many passwords can be made if repetition of letters and digits is allowed?
------------------------------------
26^4*10^2 = 45,697,600
===========================
Cheers,
Stan H.

Answer by edjones(8007)

26P4
=26!/(26-4)!
=26*25*24*23*22!/22!
=358,800
.
10P2
=10!/(10-2)!
=10*9*8!/8!
=10*9
=90
.
358800*90
=32,292,000
.
Ed